how do you solve this linear system by using substitution?
-x+2y=6
8x+y=31

Answer by mananth(16946)

-1 x + 2 y = 6 .............1
2 y = 6 + 1 x
/ 2 =
y = 3 + (1/2) x
8 x + 1 y = 31 .....................2
Substitute y in (2)
8x+1(3+(1/2)x ) = 31
8x+3+(1/2)x = 31
8x+(1/2)x = 28
8.5 x = 28
/ 8.5
x= 56/17
Plug the value of x in (1)
-1 x + 2 y = 6
-1 * 56/17 + 2 y = 6
-56/17 + 2 y = 6
2 y = 158/17
/ 2
y= 69/17
m.ananth@hotmail.ca